A cellular phone company charges $50 for each phone plus $25 per month under a special sale plan. The total cost C (in dollars) of operating the phone is given by the linear model C=50+25t where t represents the number of months of operation of a cellular phone.
How many months has the phone been in service if the total cost is $250.00?
What is the cost of the service after 18 months?
Thank You!!!

Answer by stanbon(75887)

C=50+25t where t represents the number of months of operation of a cellular phone.
How many months has the phone been in service if the total cost is $250.00?
---------------
Solve 50+25t = 250.00
25t = 200
t = 8
-----------------------------
What is the cost of the service after 18 months?
C(18) = 50+25*18 = $500
==========================
Cheers,
Stan H.